# Formalization of Auslander–Buchsbaum–Serre criterion in Lean4

### v2

@@ 1. Introduction
This project began with the goal of developing the theories of depth and Cohen–Macaulay rings. As the concepts of regular element and regular sequence were already formalized, this goal was within reach using the current library mathlib [15]. The effort was not only of mathematical interest, but also a test of the newly developed homological algebra in mathlib via its application to commutative algebra. As we shifted our focus toward combining it with homological algebra, we set Hilbert’s Syzygy theorem and the Auslander–Buchsbaum–Serre criterion as our main targets. However, at the time of this work, the $\operatorname{Tor}$ functor was still largely missing from mathlib, and the development of the Koszul complex was also incomplete, so the most well-known proof could not be used in the formalization. The turning point was finding an approach to the global dimension of a regular local ring by proving that a maximal Cohen–Macaulay module over a regular local ring is free; then the target became attainable. Later, inspired by the proof of the Ferrand–Vasconcelos theorem in [5, Theorem 2.2.8], we realized that induction on span rank (minimal number of generators) could still be carried out by focusing only on the unique maximal ideal. This allowed us to complete the full Auslander–Buchsbaum–Serre criterion. The value of this work lies not only in the final theorems: throughout the process, we formalized more basic constructions in commutative and homological algebra (for example, the definitions of projective dimension and $\operatorname{depth}$), tested these definitions in later developments, and improved their interfaces to make them easier to use. However, some of these additional interfaces are not included in this paper due to space limitations. Currently part of subsection  2.1,  2.5 and all of subsection  2.4,  2.2 are already in mathlib. We plan to upstream most of our work into mathlib in the future.
@@
If an element $x$ is both $R$-regular and $M$-regular, then $M/xM$ is a maximal Cohen–Macaulay module over $R/(x)$ by the following proposition. This allows Theorem  1.5 to be proved by induction on $\dim(R)$.
@@
$\operatorname{depth}(N,M/xM)+1=\operatorname{depth}(N,M)$.
@@
Throughout the paper, we use the symbol W for external links. Most statements and definitions are accompanied by such a link directly to the source code for the corresponding statement in mathlib or in a branch of our fork W. To keep the links stable, the links to the latter point to a fixed commit of the master branch (shortly after cleanup of the project); while other links point to the documentation of mathlib. In some code excerpts the dot notation is exhibited: for example, for I of type Ideal R, the call Ideal.depth I M can be shortened to I.depth M.
@@ 2.1. Associated Primes
theorem associatedPrimes.finite (A : Type u) [CommRing A] (M : Type v) [AddCommGroup M] [Module A M] [IsNoetherianRing A] [Module.Finite A M] : (associatedPrimes A M).Finite
@@ 2.2. Krull Dimension of Module
If $R$ is local, for a finitely generated $R$-module $M$ and an $M$-regular sequence $(r_{1},\cdots,r_{k})$, $\dim(M/(r_{1},\cdots,r_{k}))+k=\dim(M)$.
@@ 2.3. Dimension of Polynomial Rings
For local ring $(R,\mathfrak{m})$, $0pt(\mathfrak{m}[X])=0pt(\mathfrak{m})$
@@
For ring $R$, let $\mathfrak{p}$ be a prime ideal of $R[X]$, then $0pt(\mathfrak{p})\leq 0pt(\mathfrak{p}\cap R)+1$
@@ 2.4. Regular Sequence Under Flat Base Change
theorem RingTheory.Sequence.IsRegular.of_faithfullyFlat_of_isBaseChange {R S : Type*} [CommRing R] [CommRing S] [Algebra R S] {M N : Type*} [AddCommGroup M] [Module R M] [AddCommGroup N] [Module R N] [Module S N] [IsScalarTower R S N] [Module.Flat R S] {f : M →[R] N} (hf : IsBaseChange S f) {rs : List R} (reg : IsWeaklyRegular M rs) :
IsWeaklyRegular N (List.map (algebraMap R S) rs)
theorem RingTheory.Sequence.IsRegular.of_faithfullyFlat_of_isBaseChange {R S : Type*} [CommRing R] [CommRing S] [Algebra R S] {M N : Type*} [AddCommGroup M] [Module R M] [AddCommGroup N] [Module R N] [Module S N] [IsScalarTower R S N] [Module.FaithfullyFlat R S] {f : M →[R] N} (hf : IsBaseChange S f) {rs : List R} (reg : IsRegular M rs) :
IsRegular N (List.map (algebraMap R S) rs)
@@ 2.5. Projective Dimension
lemma projectiveDimension_eq_iSup_localizedModule_prime [Small.{v} R] [IsNoetherianRing R]
(M : ModuleCat.{v} R) [Module.Finite R M] :
@@ 2.6. Global Dimension
The supremum over all modules can actually be reduced to the supremum over all finitely generated modules. However, since in mathlib projective dimension is defined via vanishing of Ext and a definition of injective dimension was missing when this formalization started, we formalized only the following weakened form of [16, Lemma 18.2].W
@@
For $R$-module $M$ and $n$ a non-negative integer, if $\operatorname{Ext}_{R}^{n+1}(R/I,M)$ vanish for all $I$ ideal of $R$, then for any $R$-module NN, $\operatorname{Ext}_{R}^{n+1}(N,M)$ vanishes.
@@
lemma globalDimension_eq_sup_projectiveDimension_finite [Small.{v, u} R] : globalDimension.{v} R =
@@ 3. Definition of Depth and Related Results
rs : List R, rs.length = n (∀ r rs, r I) RingTheory.Sequence.IsRegular M rs
].TFAE
@@ 3.1. Definition of Depth
lemma moduleDepth_eq_sSup_length_regular [IsNoetherianRing R] (I : Ideal R)
(N M : ModuleCat.{v} R) [Module.Finite R M] [Nfin : Module.Finite R N]
[Nontrivial M] [Nntr : Nontrivial N] (smul_lt : I · (Top.top : Submodule R M) < Top.top)
@@
For an $M$-regular element $x\in\mathrm{Ann}(N)$, we have $\operatorname{depth}(N,M/xM)+1=\operatorname{depth}(N,M)$.
lemma moduleDepth_quotSMulTop_succ_eq_moduleDepth (N M : ModuleCat.{v} R) (x : R)
(reg : IsSMulRegular M x) (mem : x Module.annihilator R N) :
@@
$\operatorname{depth}(N,M)+k=\operatorname{depth}(N,M/(r_{1},\cdots,r_{k}))$
@@ 3.2. The Auslander–Buchsbaum Theorem
In this section, we prove the following theorem first introduced in [2].
@@
For $M$ a finitely generated free module over $R$, if the range of $f:M\rightarrow N$ is contained in $IN$ where $I$ is an ideal of $R$, then $f\in Iy\cdot(M\rightarrow N)$
@@ 4. Cohen–Macaulay Modules and Rings
For $M$ a finitely generated module over $R$ and $x$ an $M$-regular element, then $M$ is Cohen–Macaulay if and only $M/xM$ is Cohen–Macaulay.
@@
For $M$ a finitely generated module over $R$ and $(r_{1},\cdots,r_{k})$ an $M$-regular sequence, then $M$ is Cohen–Macaulay if and only $M/(r_{1},\cdots,r_{k})$ is Cohen–Macaulay.
@@
lemma isLocalize_at_prime_dim_eq_prime_depth_of_isCohenMacaulay
[Module.Finite R M] [M.IsCohenMacaulay] [ntr : Nontrivial Mₚ] :
Module.supportDim R M = p.depth M
@@
This is equivalent to (ModuleCat.of R R).IsCohenMacaulay when $R$ is a local ring.W
@@ 4.2. Universally Catenary
theorem Polynomial.isCM_of_isCM [IsNoetherianRing R] [IsCohenMacaulayRing R] :
IsCohenMacaulayRing R[X]
Assume $R$ is Cohen–Macaulay. For any prime ideal $\mathfrak{p}$ of $R[X]$, set $\mathfrak{q}=\mathfrak{p}\cap R$. Note that $R[X]\mathfrak{p}$ is the localization of $R\mathfrak{q}[X]$ at a prime ideal, so the problem reduces to the following case: if $R$ is a Cohen–Macaulay local ring and $\mathfrak{p}$ is a prime ideal of $R[X]$ with $\mathfrak{p}\cap R$ equal to the maximal ideal of $R$, then $R[X]_{\mathfrak{p}}$ is a Cohen–Macaulay local ring.W
lemma Polynomial.localization_at_comap_maximal_isCM_isCM [IsNoetherianRing R]
[IsCohenMacaulayLocalRing R] (p : Ideal R[X]) [p.IsPrime] (max : p.comap C = maximalIdeal R) :
@@ 5.1. Regular Local Ring is Domain
lemma quotient_isRegularLocalRing_tfae [IsRegularLocalRing R] (S : Finset R) (sub : (S : Set R) maximalIdeal R) :
@@
For regular local ring $R$, if $x\in\mathfrak{m}\setminus\mathfrak{m}^{2}$, then $R/(x)$ is regular local ring of dimension $\dim(R)-1$.
@@
A regular local ring is domain.
@@
For $R$ regular local ring of dimension $d$ and a sequence $rs$ of length $d$ generating its maximal ideal, any quotient of $R$ by a subsequence of $rs$ is still regular and hence a domain by the result above; therefore, $rs$ is an $R$ regular sequence.W
@@ 5.2.1. Maximal Cohen–Macaulay Module
We proceed by induction on $\dim(R)$. If $\dim(R)=0$, then $R$ is a field and $M$ is free. Assume the statement for $n$, and let $\dim(R)=n+1$. Choose $x\in\mathfrak{m}\setminus\mathfrak{m}^{2}$; by Corollary  5.2, the quotient $R/(x)$ is regular local ring of dimension $n$. To show $x$ is $M$-regular it suffices to prove that for every $\mathfrak{p}\in\mathrm{Ass}(M)$ we have $x\notin\mathfrak{p}$. By Corollary  3.13, $\dim(R/\mathfrak{p})\geq\operatorname{depth}(M)=\dim(R)=n+1$. If $x\in\mathfrak{p}$, then $n=\dim(R/(x))\geq\dim(R/\mathfrak{p})$, causing a contradiction. Hence $x$ is $M$-regular. From Theorem  3.6, we know that $\operatorname{depth}(M/xM)=\operatorname{depth}(M)-1=\dim(R)-1=n=\dim(R/(x))$, so $M/xM$ is maximal Cohen–Macaulay module over $R/(x)$. By the induction hypothesis, $M/xM$ is free over $R/(x)$, the result then follows from Lemma  5.5. ∎
@@ 5.2.2. The Main Proof
For $i=0$, $\operatorname{depth}(M)\geq\dim(R)$ implies equality, so $M$ is maximal Cohen–Macaulay and hence free by Theorem  5.7; thus $M$ has finite projective dimension. Assume the claim for $i$. For $i+1$, suppose $M\neq 0$, take a short exact sequence $0\to N\to F\to M\to 0$ with $F$ free and finitely generated. By Lemma  3.5, $\operatorname{depth}(N)+i\geq min{\operatorname{depth}(F),\operatorname{depth}(M)+1}+i=min{\operatorname{depth}(F)+i,\operatorname{depth}(M)+i+1}$. Since $\operatorname{depth}(F)=\operatorname{depth}(R)=\dim(R)$, the right-hand side is at least $\dim(R)$, so $\operatorname{depth}(N)+i\geq\dim(R)$. By the induction hypothesis, $N$ has finite projective dimension; hence, so does $M$. This completes the induction. ∎
@@ 5.3. A Weakened Version of Ferrand–Vasconcelos Theorem
lemma projectiveDimension_eq_quotient [Small.{v} R] [IsLocalRing R] [IsNoetherianRing R]
(M : ModuleCat.{v} R) [Module.Finite R M]
@@
lemma exist_isSMulRegular_of_exist_hasProjectiveDimensionLE [IsLocalRing R] [IsNoetherianRing R]
[Small.{v} R] (nebot : maximalIdeal R ⊥)
@@
Therefore $\tilde{\mathfrak{m}}$ is a direct summand of $\mathfrak{m}/x\mathfrak{m}$ as an $R/(x)$ module. By Lemma  5.9, $\operatorname{projdim}_{R/(x)}\ \mathfrak{m}/x\mathfrak{m}=\operatorname{projdim}_{R}\mathfrak{m}<\infty$, so $\tilde{\mathfrak{m}}$ has finite projective dimension. Applying the induction hypothesis, $\tilde{\mathfrak{m}}$ is generated by an $R/(x)$-regular sequence $rs$. Prepending $x$ yields $x::rs$, an $R$-regular sequence generating $\mathfrak{m}$. This completes the induction. ∎
@@
theorem IsRegularLocalRing.of_maximalIdeal_hasProjectiveDimensionLE [IsLocalRing R] [IsNoetherianRing R] [Small.{v} R]
@@
theorem IsRegularLocalRing.of_globalDimension_lt_top [IsLocalRing R] [IsNoetherianRing R]
[Small.{v} R] (h : globalDimension.{v} R < Top.top) : IsRegularLocalRing R

### v3

@@ 1. Introduction
This project began with the goal of developing the theories of depth and Cohen–Macaulay rings. As the concepts of regular element and regular sequence were already formalized, this goal was within reach using the current library mathlib [15]. The effort was not only of mathematical interest, but also a test of the newly developed homological algebra in mathlib via its application to commutative algebra. As we shifted our focus toward combining it with homological algebra, we set Hilbert’s Syzygy theorem and the Auslander–Buchsbaum–Serre criterion as our main targets. However, at the time of this work, the $\operatorname{Tor}$ functor was still largely missing from mathlib, and the development of the Koszul complex was also incomplete, so the most well-known proof could not be used in the formalization. The turning point was finding an approach to the global dimension of a regular local ring by proving that a maximal Cohen–Macaulay module over a regular local ring is free; then the target became attainable. Later, inspired by the proof of the Ferrand–Vasconcelos theorem in [5, Theorem 2.2.8], we realized that induction on span rank (minimal number of generators) could still be carried out by focusing only on the unique maximal ideal. This allowed us to complete the full Auslander–Buchsbaum–Serre criterion. The value of this work lies not only in the final theorems: throughout the process, we formalized more basic constructions in commutative and homological algebra (for example, the definitions of projective dimension and $\operatorname{depth}$), tested these definitions in later developments, and improved their interfaces to make them easier to use. However, some of these additional interfaces are not included in this paper due to space limitations. Currently part of subsection  2.5 and all of subsections  2.1,  2.2,  2.4 are already in mathlib. We plan to upstream most of our work into mathlib in the future.
@@
If an element $x\in\mathfrak{m}\setminus\mathfrak{m}^{2}$ is $M$-regular, then $M/xM$ is a maximal Cohen–Macaulay module over $R/(x)$ by the following proposition. This allows Theorem  1.5 to be proved by induction on $\dim(R)$.
@@
$\operatorname{depth}(N,M/xM)+1=\operatorname{depth}(N,M)$
@@
Throughout the paper, we use the symbol W for external links. Most statements and definitions are accompanied by such a link directly to the source code for the corresponding statement in mathlib or in a branch of our fork W. To keep the links stable, the links to the latter point to a fixed commit of the master branch; while other links point to the documentation of mathlib. In some code excerpts the dot notation is exhibited: for example, for I of type Ideal R, the call Ideal.depth I M can be shortened to I.depth M.
@@ 2.1. Associated Primes
theorem associatedPrimes.finite (A : Type u) [CommRing A] (M : Type v) [AddCommGroup M] [Module A M]
[IsNoetherianRing A] [Module.Finite A M] : (associatedPrimes A M).Finite
@@ 2.2. Krull Dimension of Module
If $R$ is local, for a finitely generated $R$-module $M$ and an $M$-regular sequence $(r_{1},\cdots,r_{k})$, $\dim(M/(r_{1},\cdots,r_{k})M)+k=\dim(M)$.
@@ 2.3. Dimension of Polynomial Rings
For a local ring $(R,\mathfrak{m})$, $0pt(\mathfrak{m}[X])=0pt(\mathfrak{m})$
@@
For a ring $R$, let $\mathfrak{p}$ be a prime ideal of $R[X]$, then $0pt(\mathfrak{p})\leq 0pt(\mathfrak{p}\cap R)+1$
@@ 2.4. Regular Sequence Under Flat Base Change
theorem IsWeaklyRegular.of_flat_of_isBaseChange {R S : Type*} [CommRing R] [CommRing S] [Algebra R S]
{M N : Type*} [AddCommGroup M] [Module R M] [AddCommGroup N] [Module R N] [Module S N]
[IsScalarTower R S N] [Module.Flat R S] {f : M →[R] N} (hf : IsBaseChange S f) {rs : List R}
(reg : IsWeaklyRegular M rs) : IsWeaklyRegular N (List.map (algebraMap R S) rs)
theorem IsRegular.of_faithfullyFlat_of_isBaseChange {R S : Type*} [CommRing R] [CommRing S] [Algebra R S]
{M N : Type*} [AddCommGroup M] [Module R M] [AddCommGroup N] [Module R N] [Module S N]
[IsScalarTower R S N] [Module.FaithfullyFlat R S] {f : M →[R] N} (hf : IsBaseChange S f) {rs : List R}
(reg : IsRegular M rs) : IsRegular N (List.map (algebraMap R S) rs)
@@ 2.5. Projective Dimension
lemma projectiveDimension_eq_iSup_localizedModule_prime
[Small.{v} R] [IsNoetherianRing R] (M : ModuleCat.{v} R) [Module.Finite R M] :
@@ 2.6. Global Dimension
The supremum over all modules can actually be reduced to the supremum over all finitely generated modules. To prove this we formalized the following form of [16, Lemma 18.2].W
@@
  4. (4)
    $\operatorname{injdim}(M)\leq n$ for all $R$-module $M$
Note that in mathlib, projective dimension and injective dimension is already defined via vanishing of $\operatorname{Ext}$; thus, this result might differ slightly from the statement in the usual sense ($(3)\leftrightarrow(4)$ is almost trivial in this setting).
@@
For $R$-module $M$ and $n$ a non-negative integer, if $\operatorname{Ext}_{R}^{n+1}(R/I,M)$ vanish for all ideals $I$ of $R$, then for any $R$-module NN, $\operatorname{Ext}_{R}^{n+1}(N,M)$ vanishes.
@@
lemma globalDimension_eq_sup_projectiveDimension_finite [Small.{v} R] : globalDimension.{v} R =
@@ 3. Definition of Depth and Related Results
rs : List R, rs.length = n (∀ r rs, r I) RingTheory.Sequence.IsRegular M rs ].TFAE
@@ 3.1. Definition of Depth
lemma moduleDepth_eq_sSup_length_regular [IsNoetherianRing R] (I : Ideal R) (N M : ModuleCat.{v} R)
[Module.Finite R M] [Nfin : Module.Finite R N] [Nontrivial M] [Nontrivial N]
(smul_lt : I · (Top.top : Submodule R M) < Top.top)
@@
For an $M$-regular element $x\in\mathrm{Ann}(N)$, we have
$\operatorname{depth}(N,M/xM)+1=\operatorname{depth}(N,M)$
lemma moduleDepth_quotSMulTop_succ_eq_moduleDepth
(N M : ModuleCat.{v} R) (x : R) (reg : IsSMulRegular M x) (mem : x Module.annihilator R N) :
@@
$\operatorname{depth}(N,M)+k=\operatorname{depth}(N,M/(r_{1},\cdots,r_{k})M)$
@@ 3.2. The Auslander–Buchsbaum Theorem
In this section, we prove the following theorem first introduced in [2].W
@@
For $M$ a finitely generated free module over $R$, if the range of $f:M\rightarrow N$ is contained in $IN$ where $I$ is an ideal of $R$, then $f\in I\cdot(M\rightarrow N)$
@@ 4. Cohen–Macaulay Modules and Rings
For $M$ a finitely generated module over $R$ and $x$ an $M$-regular element, then $M$ is Cohen–Macaulay if and only if $M/xM$ is Cohen–Macaulay.
@@
For $M$ a finitely generated module over $R$ and $(r_{1},\cdots,r_{k})$ an $M$-regular sequence, then $M$ is Cohen–Macaulay if and only if $M/(r_{1},\cdots,r_{k})M$ is Cohen–Macaulay.
@@
lemma isLocalize_at_prime_dim_eq_prime_depth_of_isCohenMacaulay [Module.Finite R M] [M.IsCohenMacaulay]
[ntr : Nontrivial Mₚ] : Module.supportDim R M = p.depth M
@@
This is equivalent to (ModuleCat.of R R).IsCohenMacaulay, but avoids passing through dimension of modules.W
@@ 4.2. Universally Catenary
theorem Polynomial.isCM_of_isCM [IsNoetherianRing R] [IsCohenMacaulayRing R] : IsCohenMacaulayRing R[X]
Assume $R$ is Cohen–Macaulay. For any prime ideal $\mathfrak{p}$ of $R[X]$, set $\mathfrak{q}=\mathfrak{p}\cap R$. Note that $R[X]_{\mathfrak{p}}$ is the localization of $R_{\mathfrak{q}}[X]$ at $\mathfrak{p}R_{\mathfrak{q}}[X]$, so the problem reduces to the following case: if $R$ is a Cohen–Macaulay local ring and $\mathfrak{p}$ is a prime ideal of $R[X]$ with $\mathfrak{p}\cap R$ equal to the maximal ideal of $R$, then $R[X]_{\mathfrak{p}}$ is a Cohen–Macaulay local ring.W
lemma Polynomial.localization_at_comap_maximal_isCM_isCM [IsNoetherianRing R] [IsCohenMacaulayLocalRing R]
(p : Ideal R[X]) [p.IsPrime] (max : p.comap C = maximalIdeal R) :
@@ 5.1. Regular Local Ring is Domain
lemma quotient_isRegularLocalRing_tfae
[IsRegularLocalRing R] (S : Finset R) (sub : (S : Set R) maximalIdeal R) :
@@
For regular local ring $R$, if $x\in\mathfrak{m}\setminus\mathfrak{m}^{2}$, then $R/(x)$ is a regular local ring of dimension $\dim(R)-1$.
@@
A regular local ring is a domain.
@@
For $R$ regular local ring of dimension $d$ and a sequence $rs$ of length $d$ generating its maximal ideal, any quotient of $R$ by a subsequence of $rs$ is still regular and hence a domain by the result above; therefore, $rs$ is an $R$-regular sequence.W
@@ 5.2.1. Maximal Cohen–Macaulay Module
We proceed by induction on $\dim(R)$. If $\dim(R)=0$, then $R$ is a field and $M$ is free. Assume the statement for $n$, and let $\dim(R)=n+1$. Choose $x\in\mathfrak{m}\setminus\mathfrak{m}^{2}$; by Corollary  5.2, the quotient $R/(x)$ is a regular local ring of dimension $n$. To show $x$ is $M$-regular it suffices to prove that for every $\mathfrak{p}\in\mathrm{Ass}(M)$ we have $x\notin\mathfrak{p}$. By Corollary  3.13, $\dim(R/\mathfrak{p})\geq\operatorname{depth}(M)=\dim(R)=n+1$. If $x\in\mathfrak{p}$, then $n=\dim(R/(x))\geq\dim(R/\mathfrak{p})$, causing a contradiction. Hence $x$ is $M$-regular. From Theorem  3.6, we know that $\operatorname{depth}(M/xM)=\operatorname{depth}(M)-1=\dim(R)-1=n=\dim(R/(x))$, so $M/xM$ is a maximal Cohen–Macaulay module over $R/(x)$. By the induction hypothesis, $M/xM$ is free over $R/(x)$, the result then follows from Lemma  5.5. ∎
@@ 5.2.2. The Main Proof
For $i=0$, $\operatorname{depth}(M)\geq\dim(R)$ implies equality, so $M$ is maximal Cohen–Macaulay and hence free by Theorem  5.7; thus $M$ has finite projective dimension. Assume the claim for $i$. For $i+1$, suppose $M\neq 0$, take a short exact sequence $0\to N\to F\to M\to 0$ with $F$ free and finitely generated. By Lemma  3.5, $\operatorname{depth}(N)+i\geq\min\{\operatorname{depth}(F),\operatorname{depth}(M)+1\}+i=\min\{\operatorname{depth}(F)+i,\operatorname{depth}(M)+i+1\}$. Since $\operatorname{depth}(F)=\operatorname{depth}(R)=\dim(R)$, the right-hand side is at least $\dim(R)$, so $\operatorname{depth}(N)+i\geq\dim(R)$. By the induction hypothesis, $N$ has finite projective dimension; hence, so does $M$. This completes the induction. ∎
@@ 5.3. A Weakened Version of Ferrand–Vasconcelos Theorem
lemma projectiveDimension_eq_quotient
[Small.{v} R] [IsLocalRing R] [IsNoetherianRing R] (M : ModuleCat.{v} R) [Module.Finite R M]
@@
lemma exist_isSMulRegular_of_exist_hasProjectiveDimensionLE
[IsLocalRing R] [IsNoetherianRing R] [Small.{v} R] (nebot : maximalIdeal R ⊥)
@@
Therefore $\tilde{\mathfrak{m}}$ is a direct summand of $\mathfrak{m}/x\mathfrak{m}$ as an $R/(x)$ module. By Lemma  5.9, $\operatorname{projdim}_{R/(x)}\ \mathfrak{m}/x\mathfrak{m}=\operatorname{projdim}_{R}\mathfrak{m}<\infty$, so $\tilde{\mathfrak{m}}$ has finite projective dimension over $R/(x)$. Applying the induction hypothesis, $\tilde{\mathfrak{m}}$ is generated by an $R/(x)$-regular sequence $rs$. Prepending $x$ yields $x::rs$, an $R$-regular sequence generating $\mathfrak{m}$. This completes the induction. ∎
@@
theorem IsRegularLocalRing.of_maximalIdeal_hasProjectiveDimensionLE [IsLocalRing R] [IsNoetherianRing R]
[Small.{v} R] (h : n, HasProjectiveDimensionLE (ModuleCat.of R (Shrink.{v} (maximalIdeal R))) n) :
@@
theorem IsRegularLocalRing.of_globalDimension_lt_top [IsLocalRing R] [IsNoetherianRing R] [Small.{v} R]
(h : globalDimension.{v} R < Top.top) : IsRegularLocalRing R